OpenStudy (anonymous):

Trig help? 1. Solve for x. sqrt 2x+3 = sqrt 6x-1 2. Solve for b. b^4/25=4

OpenStudy (anonymous):

1. \[\sqrt{2x+3} = \sqrt{6x-1}\] 2. \[\sqrt{b^4/25} = 4\]

OpenStudy (anonymous):

for the first one square both sides... you'll get 2x+3=6x-1 ... can you finish the first one now ? :)

OpenStudy (anonymous):

2x-6x=-1-3 -4x=-4 we multiply both sides by -1 4x=4 x=4/4 x=1

OpenStudy (anonymous):

is it 1??

OpenStudy (anonymous):

yes it is.. well done. For the second one I'll give you a hint... let's see what you can do :)

OpenStudy (anonymous):

\[\large \sqrt{\frac{b^4}{25}}=4\] \[\large \frac{\sqrt{b^4}}{\sqrt{25}}=4\] \[\large \frac{b^2}{5}=4\] can you do it now ? :)

OpenStudy (anonymous):

is it +/- 2 √5 ??

OpenStudy (anonymous):

yes it is... well done :)

OpenStudy (anonymous):

Thank you so much!!! :)

OpenStudy (anonymous):

My pleasure ..